OpenStudy (anonymous):

solution or no solution? 4(3m+4)=2(6m+8) 5x+2x-3=-3x+10x -(3z+4)=6z-3(3z+2) -2(j-3)=-2j+6

Nnesha (nnesha):

if you get a value of variable than yes its solution if variable cancel out then yes its no solution

OpenStudy (anonymous):

the first one is solution>>>>>>> 12m+16=12m+16 the second one is no solution>>> 7x-3=7x the third one is no solution>>>>> -3z-4=-3z+2 the fourth one is solution>>>>>> -2j+6=-2j+6

OpenStudy (anonymous):

omg thanks so much i tried but i wasnt sure if i was doing it right! :)

OpenStudy (anonymous):

no problem!

OpenStudy (radar):

First one is an identity, with an infinite number of solutions.

OpenStudy (radar):

Second one is inconsistent with no solution.

OpenStudy (radar):

Third one is inconsistent with no solution.

OpenStudy (radar):

The last one is an identity with infinite number of solutions.

OpenStudy (radar):

Just test them out you will discover the problems.